OpenStudy (anonymous):

write the slope-intercept forms of the equations of the lines through the given point (a) parallel to the given line and (b) perpendicular to the given line. Point (3,-2) and the line is 5x-4y= 8

OpenStudy (jack1):

Point (3,-2) and the line is 5x-4y= 8 slope-intercept form y = mx +c so 5x-4y= 8 5x- 8= 4y y = (5/4) x - 2 ==> slope intercept form of original line

OpenStudy (jack1):

a) parallel to this line y = (5/4) x + c and passes through 3,-2 -2 = (5/4) 3 + c -2 = 15/4 + c c = -2 - 15/4 = -2 -3.75 = -5.75 so equation of the parallel line is y = (5/4) x - 5.75

OpenStudy (goformit100):

@Jack1 is correct

OpenStudy (jack1):

@goformit100 ... dude, you've blocked me... i cant pm u... why?